Dyson Company in Senai Johor plans to produce a new hair product known as Dyson Supersonic. The suppliers for the company are Company A and Company B found that 1% and 2% defective of 200 and 300 items respectively. Arrange appropriate hypothesis testing to investigate whether Company B is better by using a 0.05 level of significance.
1
Expert's answer
2019-08-06T10:35:25-0400

Null hypothesis "H_0:\;p_1=p_2"

Alternative hypothesis "H_a:\;p_1< p_2"

Pooled sample proportion: "E=mc^2p=\frac{p_1n_1+p_2n_2}{n_1+n_1}=\frac{0.01*200+0.02*300}{200+300}=0.016"

Test statistic: "z=\frac{p_1-p_2}{\sqrt{p(1-p)(\frac{1}{n_1}+\frac{1}{n_2})}}=-0.87."

P-value: "p=0.191."

Since the P-value is greater than 0.05, reject the null hypothesis.